pls. help me solve this prob. thank you....
1. find the equation of the line through (-3,2) with slope -1/3

2. find the equation of the line parallel to 3x-2y=3 and passing through (4,6).
3. find the equation of the line through the y-intercepts of the lines 3x+4y-12=0 and 3x-y-9=0.

Answer by ewatrrr(10682)


Hi
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
1. y = (-1/3)x + b |using ordered pair Pt(-3,2) to solve for b
2 = (-1/3)*-3 + b
1 = b
y = (-1/3)x + 1
2. parallel line have the same slope: 3x-2y=3 OR y = (3/2)x - 3/2 m = (3/2)
y = (3/2)x + b |using ordered pair Pt(4,6) to solve for b
6 = (3/2)*4 + b
0 = b
y = (3/2)x
3. 3x+4y-12=0 OR y = (-3/4)x + 3 y-intercept is Pt(0,3)
3x-y-9=0 OR y = 3x - 9 y-intercept is Pt(0,-9)
Line thru these y-intercepts is x = 0